OpenStudy (anonymous):

Achieve $225,500 at 8.15% compounded continuously for 8 years, 155 days.

OpenStudy (ageta):

8 years, 155 days = 3075/365 years, assuming 1 year = 365 days F = Pe^(rt) 225,500 = Pe^(0.0815*3075/365) P = 225,500/e^(0.0815*3075/365) = $113,489.5282

OpenStudy (ageta):

do you understand?